OpenStudy (anonymous):

x - 2y = 4 3x + y = 6 How do I solve this and graph it out?

OpenStudy (anonymous):

x-2y=4 -2y=4-x y=(4-x)/-2 3x+y=6 y=6-3x plug in values for x like 1, 2, 3 which would be the X coordinate then solve to find what Y equals and that give you coordinates do it multiple time to get many points